### v1

@@ 4.2 Object Detection and Instance Segmentation
Results: The experimental results on the COCO dataset are detailed in Tables 3 and 4 by using RetinaNet and Mask R-CNN respectively. Apparently, SAEViT ranks first in each group. Especially, SAEViT-T and SAEViT-XS achieve 39.5 AP and 41.8 AP respectively, surpassing other competitors with safe margins. Similarly, using Mask R-CNN detector, SAEViT-T and SAEViT-XS achieve 40.1 AP and 42.6 AP respectively, which exhibits strong object localization and segmentation ability under complex scenes.
@@
Visualizations: Fig. 4 shows detection results by using different detectors. The $1^{st}$ and the $3^{rd}$ columns shows the results of SAEViT-XS by using RetinaNet, the $2^{nd}$ and the $4^{th}$ columns show the results by using Mask R-CNN. Generally, they have all achieved detection tasks for commonly known objects, except that RetinaNet has some missed detections, while Mask R-CNN made some duplicated detections.
@@ 4.3 Semantic Segmentation on ADE20K
Results:On the ADE20K dataset, the results of SAEViT in semantic segmentation tasks are presented in Table 5. SAEViT-T achieves 40.6 mIoU, showing improved segmentation accuracy across different semantic categories and demonstrating precise recognition and segmentation of various objects and scenes within images. Meanwhile, SAEViT-XS obtains 42.1 mIoU, ensuring both accuracy and integrity in semantic segmentation under complex scenes.
@@ 4.4 GPU Throughput and Latency tests
As the key indicator for measuring the real-time inference performance of deep learning models, while device latency directly reflects the response speed of the model in practical application scenarios such as real-time image recognition and autonomous driving decision systems. In this section, throughput and latency tests are implemented multiple times to measure the model’s average throughput (i.e. the number of images processed per second) and single-image inference latency (milliseconds) on a given device. The experimental setup includes a 5s warm-up period, random initialization of 224×224 feature maps, and batch sizes set to 256 for throughput measurement and 1 for latency measurement; to ensure fairness and reduce error, forward propagation is run 50 times and the delay averaged. All tests are conducted on the same AMD Ryzen 7 5800H CPU and NVIDIA RTX 3060 Laptop GPU, comparing PVT-T, ShuntedViT-T, CMT-Ti and our model SAEViT-XS.
The results are shown in Table 6. Our model has the minimum number of parameters and relatively low computation cost. However, we achieve slightly lower accuracy than ShuntedViT-T, we realize the highest throughput and lowest latency among all the methods by safe margins.
@@ 4.5.4 The role of the conv-stem
The conv-stem plays a critical role in ViT families. We propose to employ a conv-stem head (instead of using a patch embedding) for down-sampling. When the input feature map is 224×224, the patch size is set to 4 in the initial stage. After the first down-sampling layer, the image dimension is reduced to 1/4 of the original size (i.e. 56×56). However, it may cause the loss of more fine-grained features. An experiment has been conducted on ImageNet-1K to testify the function of conv-stem on accuracy. As shown in Table 9, using conv-stem brings obvious higher accuracy without substantially increasing computational costs.
@@ 5 Conclusion
In this paper, we propose SAEViT, a hybrid ViT based architecture for various vision tasks. The core components include Sparsely Aggregated Attention and channel-interactive Feed-Forward Network. By fusing the local prior capability of the convolution blocks with ViT’s global modeling strengths, SAEViT significantly enhances the performances while maintaining relatively low computational overhead. Extensive experiments have been conducted on 3 different vision tasks to validate its effectiveness, where the advantage in the accuracy-speed tradeoff has been highlighted against other methods. Additional ablation experiments are also carried out to verify the function of the core components. Although effective, SAEViT has some limitations under special scenarios. Firstly, the SAA module performs global static downsampling based on token redundancy, but for areas with rich details or small targets, it may cause excessive compressed information and therefore results in lower recognition accuracy for fine-grained features. Meanwhile, its fixed sampling rate also limits its adaptability to high-resolution inputs. Secondly, although CIFFN enhances inter-channel feature interaction, it still induces certain computational and memory overhead due to the depth-wise separable convolution and re-assignment operations, which may affect deployment efficiency in resource constrained environments. In the future, we will focus on developing dynamic or layered sparse sampling methods to preserve key information, and introducing learnable multi-scale attention mechanisms to enhance adaptability to different input resolutions. And in the future, we will develop more efficient sparse sampling strategies and deploy the model to edge devices to exert its role in practical situations.

### v2

@@ 4.2 Object Detection and Instance Segmentation
Results: The experimental results on the COCO dataset are detailed in Tables 3 and 4 by using RetinaNet and Mask R-CNN respectively. Apparently, SAEViT ranks first in each group. Especially, SAEViT-T and SAEViT-XS achieve 39.5 AP and 41.8 AP respectively, surpassing other competitors with safe margins. Similarly, using Mask R-CNN detector, SAEViT-T and SAEViT-XS achieve 40.1 AP and 42.6 AP respectively, which exhibits strong object localization and segmentation ability under complex scenes. PoolFormer-S12[35], PVTv2-B0[33] and EdgeViT-XS[23] only achieve 36.2 AP, 37.2 AP and 38.7AP respectively. EdgeViT-XS adopts a local-global-local delegate-token mechanism, which reduces the cost of self-attention while performing strong aggregation and sparsification operations on spatial information. In contrast, our SAA uses learnable upsampling (deconvolution + skip connections) to recover more high-frequency details, which makes it more advantageous in detecting small object. On the other hand, CIFFN enhances inter-channel interaction, allowing the model to better discriminate object features from cluttered backgrounds. The complementary function of SAA+CIFFN forms a strong backbone, providing richer and more discriminative multi-scale feature representations. Similarly, SAEViT-XS achieves 41.8 AP with RetinaNet, surpassing SBCFormer-L[19](41.1 AP) and EdgeViT-XS[23](40.6 AP). The improvement is more prominent by using Mask R-CNN, where SAEViT-XS reaches 42.6 box AP and 43.5 mask AP, outperforming RepViT-M1.1[30](39.8 AP) and EfficientFormer-L3[13](41.4 AP). These results confirm the superiority of our multi-scale hierarchical design in enhancing feature representation for both object localization and segmentation. The ability to preserve spatial details through SAA and enrich channel interactions via CIFFN allows our model to excel in complex scenes with overlapping objects and varied scales.
@@
Visualizations: Fig. 4 shows detection results by using different detectors. The $1^{st}$ and the $3^{rd}$ columns shows the results of SAEViT-XS by using RetinaNet, the $2^{nd}$ and the $4^{th}$ columns show the results by using Mask R-CNN. Generally, SAEViT-based detectors maintain high precision and recall, which have successfully identified most commonly known objects(except that RetinaNet has some missed detections, while Mask R-CNN made some duplicated detections). Since our model preserves the fine-grained details through sparse attention and enhances channel interactions, which greatly improve the detection results.
@@ 4.3 Semantic Segmentation on ADE20K
Results: As shown in Table 5, SAEViT achieves appealing performance on ADE20K semantic segmentation benchmarks. SAEViT-T achieves 40.6 mIoU with only 10.1M parameters, while SAEViT-XS attains 42.1 mIoU with 13.1M parameters, outperforming PVTv2[33], PoolFormer[35], and EdgeViT [23] under the same level of parameter constraints. PVTv2 emphasizes hierarchical token processing and PoolFormer focuses on a minimalist token-mixer, which all ignore the recovery and fusion of multi-scale details with reduced dimensionality. In contrast, SAEViT has strong capability in capturing and reconstructing semantic contexts across diverse scenes. The performance improvement can be largely attributed to the model’s multi-scale feature aggregation mechanism enabled by SAA and enhanced inter-channel communication through CIFFN. Considering that precise pixel-level classification and context integration are critical for segmentation tasks, SAA preserves structural details even after downsampling via learnable deconvolution, maintaining boundary information. Meanwhile, CIFFN promotes more discriminative feature learning through channel-wise re-assginment, effectively reducing intra-class ambiguity and ensuring segmentation consistency, especially for categories with high visual similarity. Furthermore, the hierarchical pyramid design of SAEViT allows the model to capture features at multiple scales, making it particularly suited for segmenting objects of varying sizes within complex scenes.
@@ 4.4 GPU Throughput and Latency tests
As the key indicator for measuring the real-time inference performance of deep learning models, while device latency directly reflects the response speed of the model in practical application scenarios such as real-time image recognition and autonomous driving decision systems. In this section, throughput and latency tests are implemented multiple times to measure the model’s average throughput (i.e. the number of images processed per second) and single-image inference latency (milliseconds) on a given device. The experimental setup includes a 5s warm-up period, random initialization of 224×224 feature maps, and batch sizes are set to 256 for throughput measurement and 1 for latency measurement. The forward propagation is performed 50 times to ensure fair comparison. All tests are conducted on the same AMD Ryzen 7 5800H CPU and NVIDIA RTX 3060 Laptop GPU, comparing PVT-T, ShuntedViT-T, CMT-Ti and our model SAEViT-XS.
Table 6 compares the throughput and latency of SAEViT-XS against other lightweight vision Transformers. Among them, SAEViT-XS achieves the highest throughput and the lowest latency on an NVIDIA RTX 3060 GPU with fewer parameters. This efficiency stems from our sparsely aggregated attention mechanism, which reduces redundant computations, and the optimized CIFFN module minimizes inter-channel redundancy without extensive linear projections. As a result, SAEViT has high suitability for real-time applications on edge devices, where both accuracy and inference speed are critical.
@@ 4.5.4 The role of the conv-stem
As shown in Table 9, incorporating a convolutional stem improves SAEViT-T’s accuracy from 74.89% to 76.28% with only a minimal increase in parameters (0.06M). This improvement proves that convolutional stems better preserve low-level features such as edges and textures, which are often lost in patch-based embeddings. The conv-stem thus serves as a lightweight yet effective feature enhancer, which is particularly beneficial for extraction of fine-grained details.
@@ 5 Conclusion
In this paper, we propose SAEViT, a hybrid ViT based architecture for various vision tasks. The core components include Sparsely Aggregated Attention and channel-interactive Feed-Forward Network. By fusing the local prior capability of the convolution blocks with ViT’s global modeling strengths, SAEViT significantly enhances the performances while maintaining relatively low computational overhead. Extensive experiments have been conducted on 3 different vision tasks to validate its effectiveness, where the advantage in the accuracy-speed tradeoff has been highlighted against other methods. Additional ablation experiments are also carried out to verify the function of the core components.
Although effective, SAEViT has some limitations under special scenarios. Firstly, the SAA module performs global static downsampling based on token redundancy, but for areas with rich details or small targets, it may cause excessive compressed information and therefore results in lower recognition accuracy for fine-grained features. Meanwhile, its fixed sampling rate also limits its adaptability to high-resolution inputs. Secondly, although CIFFN enhances inter-channel feature interaction, it still induces certain computational and memory overhead due to the depth-wise separable convolution and re-assignment operations, which may affect deployment efficiency in resource constrained environments.
In the future, we will be focusing on 2 research directions. First, we will investigate a dynamic sparse attention mechanism that adaptively adjusts the sampling rate based on image content. For instance, a lightweight prediction module could dynamically allocate sampling rates according to local texture complexity (i.e. assign a smaller sr to texture-rich regions to preserve details and a larger sr for smooth areas to enhance computational efficiency). Second, we will conduct extensive evaluation of real-device deployment and energy efficiency, which aims to test the practical FLOPs on certain edge devices. To be specific, we plan to measure single-batch inference latency, peak memory usage, and energy consumption (in Joules per inference) across a diverse set of hardware platforms. A comparative analysis will be performed against state-of-the-art lightweight models. Key evaluation metrics will include latency (ms), frames per second (FPS), power consumption (J/inference), as well as Top-1 accuracy and mAP after quantization. These experiments will provide critical insights into the practical applicability of SAEViT in real-world edge computing scenarios. We hope our research outcome could pave the way for future researchers in this direction.